OpenStudy (anonymous):

Find the derivative of the function f(x)=(2x-3)^4 (x^2 +x+1)^5 Im not sure I am doing this correctly. I started off using the product rule and got (x^2+x+1)^5 d/dx(2x-3)^4 + (2x-3)^4 d/dx(x^2+x+1)^5 I then used the chain rule on the d/dx terms leaving me with (x^2+x+1)^5 4(2x-3)^3 *2 + (2x-3)^4 5(x^2+x+1)^4 *2x+1 I seem to be far off from the answer. The answer in the book is (2x-3)^3 (x^2+x+1)^4 (28x^2-12x-7) I have no clue how to get that answer. Help would be GREATLY APPRECIATED! Thanks

OpenStudy (mandre):

You have to start with the product rule then use the chain rule as you have derivatives of composite functions.

OpenStudy (mandre):

\[\frac{ d }{ dx }(x^2+x+1)^5 = 5(x^2+x+1)^4.(2x + 1)\]